# Exploring Context-aware and LLM-driven Locomotion for Immersive Virtual Reality

### v1

@@ Abstract
Locomotion plays a crucial role in shaping the user experience within virtual reality environments. In particular, hands-free locomotion offers a valuable alternative by supporting accessibility and freeing users from reliance on handheld controllers. To this end, traditional speech-based methods often depend on rigid command sets, limiting the naturalness and flexibility of interaction. In this study, we propose a novel locomotion technique powered by large language models (LLMs), which allows users to navigate virtual environments using natural language with contextual awareness. We evaluate three locomotion methods: controller-based teleportation, voice-based steering, and our language model-driven approach. Our evaluation measures include eye-tracking data analysis, including explainable machine learning through SHAP analysis as well as standardized questionnaires for usability, presence, cybersickness, and cognitive load to examine user attention and engagement. Our findings indicate that the LLM-driven locomotion possesses comparable usability, presence, and cybersickness scores to established methods like teleportation, demonstrating its novel potential as a comfortable, natural language-based, hands-free alternative. In addition, it enhances user attention within the virtual environment, suggesting greater engagement. Complementary to these findings, SHAP analysis revealed that fixation, saccade, and pupil-related features vary across techniques, indicating distinct patterns of visual attention and cognitive processing. Overall, we state that our method can facilitate hands-free locomotion in virtual spaces, especially in supporting accessibility.
@@
With the rapid advancement of virtual reality (VR) technologies, VR systems have started to be used widely in various domains and purposes, such as education [33], entertainment [5], healthcare [35] and training [79]. The quality of user interaction within immersive environments is essential to shape user experience positively. One of the important aspects of user experience is locomotion, which refers to how users move within virtual environments. Locomotion affects many factors including user immersion [12], task performance [26], overall comfort [39], and more importantly the cybersickness [25]. Among various techniques, teleportation by using handheld controllers remains one of the most widely adopted methods [62].
@@
Recent advancements in large language models (LLMs) such as GPT-4 offer new opportunities for developing more natural and intelligent voice-based interaction systems in VR [15]. These models can accurately interpret user instructions expressed in natural language in VR and effectively identify intent, particularly when the environmental context is incorporated into the prompt. Moreover, as they support multiple languages, it is possible to enable multilingual interaction within the same system setup and make virtual environments more generic with minimal effort while supporting user diversity.
@@ 2 Methodology
In this study, we investigated the effectiveness of three distinct locomotion techniques in VR: teleportation, voice-based steering through fixed commands [39], and our proposed LLM-driven locomotion approach. To evaluate these techniques, we designed a virtual environment where participants were given a task that required them to navigate through the scene in our between-subjects design user study. We collected both objective data, including eye tracking to assess attention and cognition, and subjective data, including the System Usability Scale (SUS), the Igroup Presence Questionnaire (IPQ), the Cybersickness in Virtual Reality Questionnaire (CSQ-VR), and the NASA Task Load Index (NASA-TLX). These questionnaires measured usability, sense of presence, cybersickness, and cognitive load. In the following subsections, we describe the locomotion techniques and explain the details of our user study. The Institutional Review Board of [Blind Institute] approved our study protocols and data collection procedures.
@@ 2.1 Locomotion Techniques
In this study, we implemented three different locomotion techniques, including teleportation, voice-based steering through fixed commands, and proposed LLM-driven locomotion. We treat the teleportation and voice-based steering as general and hands-free baselines, respectively.
@@ 2.1.1 Controller-based Teleportation
We use controller-based teleportation as the baseline locomotion method, representing a conventional and widely adopted technique in VR [11, 3]. This method is typically perceived as easy to learn and use, and we chose it because of its easy adaptability to most virtual environments. Users are generally comfortable with this locomotion technique [62]. In this method, the user points to a location using the controller button and is instantly transported to the selected position. In the default implementation, a curved line is displayed to guide the user. A green arc indicates a valid teleportation target, while a red arc signals that teleportation to the selected area is not possible. When a valid location is selected, the user is immediately transported to this location after participants pull their fingers from the button. This method minimizes motion-induced discomfort and is generally well-tolerated by users [11].
@@ 2.1.2 Voice-based Steering through Fixed Commands
Although it is mainly designed as a hands-free alternative for our experiment setting to minimize any errors or effects of design preferences, we use a trigger button to start talking and use the same button to stop recording to minimize the failures in voice recognition. Once a valid command such as <go forward> is provided, the user starts moving continuously at a constant speed. While directional changes could be controlled using the <turn left> and <turn right> commands, users are also free to rotate naturally using their physical head and body movements. Nevertheless, unlike teleportation, which involves instantaneous movement and minimizes sensory conflict, voice-based steering relies on continuous movement, which has been shown to increase the likelihood of VR-induced motion sickness [39]. While teleportation kind instant movement-based methods do not trigger expectations of physical motion, continuous movement generally introduces a greater mismatch between visual and vestibular signals, leading to increased discomfort [55].
@@ 2.1.3 LLM-driven Locomotion
To support this, we utilized the pipeline illustrated in Figure Exploring Context-aware and LLM-driven Locomotion for Immersive Virtual Reality, which takes the system prompt, the user’s voice input, and contextual scene information to construct a comprehensive user prompt. The system prompt primarily describes the purpose of the agent as a “navigation assistant in Unity” and explains the structure of the user prompts, specifying the rules and expected output format. It also provides three examples. The user prompt includes the user’s command and contains explicit environmental context, including the positions of nearby landmarks, buildings, and vehicles currently visible to the user. The LLM processes this information to generate precise target coordinates. Subsequently, these coordinates are dynamically mapped to the closest valid navigable location within the virtual environment based on predefined walkable areas. In scenarios where the generated coordinates correspond to multiple possible valid locations, such as intersections or corners, the system employs the user’s gaze direction to select the most appropriate position by calculating the smallest angle between the gaze vector and available street directions to align with the user’s intended movement trajectory. Our system primarily relies on eye gaze for this calculation; however, head gaze is used as a fallback if eye gaze data is unavailable or unreliable. Additionally, to ensure robustness, the system incorporates several fallback mechanisms. If the LLM returns excessive text along with the location rather than only the expected coordinate, we dynamically extract the coordinate from the response. If the LLM fails to generate meaningful or interpretable output, or if the user’s command is ambiguous, the user’s position remains unchanged, effectively preventing unintended movement. Successful commands are visually indicated to the user by a curved line marking the calculated destination point, and after a two-second interval, the user is transported to the new location. Similar to the voice-based steering through fixed commands, we use a trigger button to start and stop voice recording during the experiment to minimize the effect of design preferences.
@@ 2.2.3 Experimental Design
We employed a between-subjects design, where each participant experienced only one of the three locomotion methods: teleportation, fixed voice command-based locomotion, or LLM-driven locomotion. This approach ensured that individual learning effects, fatigue, or cybersickness did not influence performance across different conditions. All participants navigated the same virtual environment. We conducted the study in a VR environment developed using Unity, designed to resemble a virtual small town with a simple street layout consisting of four vertical and four horizontal roads, as shown in Figure 1. We designed a two-step navigation task to evaluate each locomotion method within a task-oriented context. We first instructed participants to locate a bank in the virtual town. Each participant started at the same position, marked by the blue “X” in Figure 1, and completed the task by finding the target locations at identical coordinates, indicated by pink holograms on the street. To make the task intuitive, we placed a prominent purple hologram in front of the bank, making it easy to identify. Upon reaching the bank, the first hologram disappeared, and a second one appeared at the town’s exit as a second target. Participants were then required to find and reach this exit point to complete the navigation task.
@@ 2.2.4 Procedure
We first welcomed the participants and asked them to provide written consent to participate in the experiment. Next, we explained the (randomly) assigned locomotion condition to them in detail. Participants then wore the HMD, and the session began with a demo scene designed to help them become familiar with the locomotion technique and VR. Eye-tracking calibration was also performed during this phase. Once participants felt comfortable about the VR and controls, following the demo, the main experiment commenced. As instructed, participants were supposed to navigate the virtual environment and reach two holograms placed at specific locations. After completing the task, participants completed questionnaires, including the NASA-TLX, SUS, IPQ, CSQ-VR, and a demographic survey. Later, they were compensated with €7.5 for an experimental session that was expected to last half an hour in total.
@@ 2.2.5 Measurements
To evaluate the locomotion techniques, we conducted an eye-tracking analysis, as eye tracking has been widely used to assess user engagement, attention, behavior, and cognitive processes during interaction [38, 80]. We classified samples as either fixations or saccades from the raw eye-tracking data. Fixations represent moments when the eyes remain steadily focused on a specific point, while saccades are rapid eye movements between fixation points. We analyzed several features to understand users’ visual attention and cognitive states. The fixation rate, which reflects the frequency of attentional shifts, is commonly used to assess how users allocate their attention during a task [28]. Mean fixation duration, which represents the length of time users dwell on specific points, has been linked to deeper cognitive processing and enhanced information encoding [43]. The mean saccade duration captures the speed of gaze transitions, indicating the efficiency of visual search behavior. Pupil diameter provides insights into the mental effort and arousal, correlating with the cognitive workload, fatigue, and engagement [6, 40]. Additionally, we identified blinks and extracted both blink frequency and duration, as these factors are associated with cognitive load and visual fatigue [7]
@@ 2.2.7 Analysis
We conducted a series of statistical analyses on both questionnaire responses (e.g., SUS, NASA-TLX) and eye-tracking measures to examine differences across experimental conditions. For each dependent variable, we performed a one-way Analysis of Variance (ANOVA) to evaluate the effect of the locomotion technique (teleportation, LLM-driven locomotion, and voice-based steering). When the ANOVA indicated a statistically significant effect, we used Tukey’s Honest Significant Difference (HSD) post-hoc test to determine which pairs of conditions differed significantly. This approach controlled the family-wise error rate while comparing all condition pairs.
We checked the assumptions of normality and homogeneity of variance before conducting ANOVA, applying Shapiro-Wilk tests for normality and Levene’s test for homogeneity of variances. In cases where these assumptions were violated, we conducted the non-parametric Kruskal-Wallis H test as an alternative to ANOVA. We conducted all statistical tests using . Where relevant, we also report the effect sizes (e.g., ) to support the interpretation of findings.
@@ 2.2.8 Model Building
We conducted detailed analyses of eye-tracking data to investigate how different locomotion techniques influence user behavior. These analyses aim to reveal potential cognitive and attentional differences among participants using teleportation, voice-based steering locomotion, and LLM-driven locomotion techniques. Additionally, we trained classification models to predict the locomotion condition based on eye-tracking features. This process enables us to identify which gaze-related features are most affected by the locomotion technique and which features are the most predictive or distinctive in differentiating between the locomotion techniques. We provide the full list of features used in Table 2.
@@
We evaluated the performance of five classifiers: Random Forest, LightGBM, k-Nearest Neighbor (k-NN), Support Vector Machine (SVM), and Gradient Boosting. The classification was performed on data from 20-second windows, with 80% of the data used for training and 20% for testing. Hyperparameters were optimized using a stratified 5-fold cross-validation strategy to ensure robust model performance across classes. To understand the models’ decision-making processes, we applied SHAP (SHapley Additive exPlanations) [50] to interpret feature importance in a transparent and model-agnostic fashion. SHAP values provide a detailed attribution of feature contributions toward classification outcomes, helping us identify which specific aspects of gaze behavior, such as fixation duration, saccade amplitude, or pupil diameter, are most influenced by the underlying locomotion technique.
@@ 3.2 Eye-Tracking Analysis
To assess cognitive engagement under different locomotion techniques, we analyzed the fixation rate, defined as the number of fixations per second, as shown in Figure 6. ANOVA revealed a statistically significant effect of locomotion type on fixation rate, , , with a large effect size (), indicating that the type of locomotion significantly influenced visual attention patterns. Additionally, the statistics indicated that the voice-based steering condition resulted in the highest fixation rate (, ), followed by teleportation (, ), and voice LLM (, ). Post-hoc comparisons using Tukey’s Honest Significant Difference (HSD) test showed that both teleportation and voice LLM conditions had significantly lower fixation rates compared to the voice-based steering condition ( for both comparisons). However, no statistically significant difference was observed between the teleportation and voice LLM conditions (), suggesting similar visual engagement across these two hands-free techniques.
@@
Mean saccade durations for each condition are presented in Figure 6, and all values are reported in milliseconds. ANOVA revealed a statistically significant effect of locomotion type on mean saccade duration, , , with a large effect size (). Descriptive statistics showed that the voice-based steering condition had the highest mean saccade duration (, ms), followed by teleportation (, ms), and the LLM-driven condition with the lowest duration (, ms). Post-hoc tests showed mean saccade duration was significantly lower in the LLM-driven condition than in teleportation () and voice-based steering (), with no difference between the latter two ().

### v2

@@ Abstract
Locomotion plays a crucial role in shaping the user experience within virtual reality environments. In particular, hands-free locomotion offers a valuable alternative by supporting accessibility and freeing users from reliance on handheld controllers. To this end, traditional speech-based methods often depend on rigid command sets, limiting the naturalness and flexibility of interaction. In this study, we propose a novel locomotion technique powered by large language models (LLMs), which allows users to navigate virtual environments using natural language with contextual awareness. We evaluate three locomotion methods: controller-based teleportation, voice-based steering, and our language model-driven approach. Our evaluation combines eye-tracking data analysis, including exploratory explainable machine learning analysis with SHAP, and standardized questionnaires (SUS, IPQ, CSQ-VR, NASA-TLX) to examine user experience through both objective gaze-based measures and subjective self-reports of usability, presence, cybersickness, and cognitive load. Our findings show no statistically significant differences in usability, presence, or cybersickness between LLM-driven locomotion and established methods such as teleportation, suggesting its potential as a viable, natural language-based, hands-free alternative. In addition, eye-tracking analysis revealed patterns suggesting tendency toward increased user attention and engagement in the LLM-driven condition. Complementary to these findings, exploratory SHAP analysis revealed that fixation, saccade, and pupil-related features vary across techniques, indicating distinct patterns of visual attention and cognitive processing. Overall, we state that our method can facilitate hands-free locomotion in virtual spaces, especially in supporting accessibility.
@@
With the rapid advancement of virtual reality (VR) technologies, VR systems have started to be used widely in various domains and purposes, such as education [33], entertainment [5], healthcare [35] and training [79]. The quality of user interaction within immersive environments is essential to shape user experience positively. An important aspect of user experience is locomotion, which refers to how users move within virtual environments. Locomotion affects many factors including user immersion [12], task performance [26], overall comfort [39], and more importantly the cybersickness [25]. Among various techniques, teleportation by using handheld controllers remains one of the most widely adopted methods [62].
@@
Recent advancements in large language models (LLMs) such as GPT-4 offer new opportunities for developing more natural and intelligent voice-based interaction systems in VR [15]. These models can interpret natural language instructions in VR and effectively identify intent, particularly when environmental context is included in the prompt. Moreover, as they support multiple languages, it is possible to enable multilingual interaction within the same system setup and make virtual environments more generic with minimal effort while supporting user diversity.
@@ 2 Methodology
In this study, we investigated the effectiveness of three distinct locomotion techniques in VR: teleportation, voice-based steering through fixed commands [39], and our proposed LLM-driven locomotion approach. To evaluate these techniques, we designed a virtual environment where participants were given a task that required them to navigate through the scene in our between-subjects design user study. We collected both objective data, including eye tracking to assess attention and cognition, and subjective data, including the System Usability Scale (SUS), the Igroup Presence Questionnaire (IPQ), the Cybersickness in Virtual Reality Questionnaire (CSQ-VR), and the NASA Task Load Index (NASA-TLX). These questionnaires measured usability, sense of presence, cybersickness, and cognitive load. In the following subsections, we describe the locomotion techniques and explain the details of our user study. The Institutional Review Board of the Technical University of Munich approved our study protocols and data collection procedures.
@@ 2.1 Locomotion Techniques
We evaluate controller teleportation, fixed-command voice steering, and our LLM-driven, context-aware teleportation to represent different locomotion approaches: teleportation as the standard baseline for efficiency and comfort, voice steering as a hands-free continuous-motion baseline using the same speech channel, and the LLM method as a hands-free, selection-based alternative using free-form, context-aware language.
@@ 2.1.1 Controller-based Teleportation
We use controller-based teleportation as the baseline locomotion method, representing a conventional and widely adopted technique in VR [11, 3]. This method is typically perceived as easy to learn and use, and we chose it because of its easy adaptability to most virtual environments. Users are generally comfortable with this locomotion technique [62]. In this method, the user points to a location using the controller button and is instantly teleported to the selected position. In the default implementation, a curved line is displayed to guide the user. A green arc indicates a valid teleportation target, while a red arc signals that teleportation to the selected area is not possible. When a valid location is selected, the user is immediately teleported to this location after participants pull their fingers from the button. This method minimizes motion-induced discomfort and is generally well-tolerated by users [11].
@@ 2.1.2 Voice-based Steering through Fixed Commands
Although it is mainly designed as a hands-free alternative for our experiment setting to minimize any errors or effects of design preferences, we use a trigger button to start talking and use the same button to stop recording to minimize the failures in voice recognition. Once a valid command such as <go forward> is provided, the user starts moving continuously at a constant speed. While directional changes could be controlled using the <turn left> and <turn right> commands, users are also free to rotate naturally using their physical head and body movements. Nevertheless, unlike teleportation, which involves instantaneous movement and minimizes sensory conflict, voice-based steering relies on continuous movement. This type of movement can create a mismatch between visual and vestibular signals, increasing the likelihood of VR-induced motion sickness [39, 55].
@@ 2.1.3 LLM-driven Locomotion
To support this, we utilized the pipeline illustrated in Figure Exploring Context-aware and LLM-driven Locomotion for Immersive Virtual Reality, which takes the system prompt, the user’s voice input, and contextual scene information to construct a comprehensive user prompt. The system prompt primarily describes the purpose of the agent as a “navigation assistant in Unity” and explains the structure of the user prompts, specifying the rules and expected output format. It also provides three examples. The user prompt includes the user’s command and contains explicit environmental context, including the positions of nearby landmarks, buildings, and vehicles currently visible to the user. This contextual information is obtained directly from objects that are both visible and have a specific tag, with their names extracted using a custom script in Unity. The LLM processes this information to generate precise target coordinates. Subsequently, these coordinates are dynamically mapped to the closest valid navigable location within the virtual environment based on predefined walkable areas. When generated coordinates correspond to multiple locations, such as intersections or corners, the system selects the street orientation that makes the smallest angle with the user’s head orientation to best match the intended movement trajectory. Additionally, to ensure robustness, the system incorporates several fallback mechanisms. If the LLM returns excessive text along with the location rather than only the expected coordinate, we dynamically extract the coordinate from the response. If the LLM fails to generate meaningful or interpretable output, or if the user’s command is ambiguous, the user’s position remains unchanged, effectively preventing unintended movement. Successful commands are visually indicated to the user by a curved line marking the calculated destination point, and after a two-second interval, the user is teleported to the new location. Similar to the voice-based steering through fixed commands, we use a trigger button to start and stop voice recording during the experiment to minimize the effect of design preferences.
@@ 2.2.3 Experimental Design
We employed a between-subjects design, where each participant experienced only one of the three locomotion methods: teleportation, fixed voice command-based locomotion, or LLM-driven locomotion. This approach ensured that individual learning effects, fatigue, or cybersickness did not influence performance across different conditions. All participants navigated the same virtual environment. We conducted the study in a VR environment developed using Unity, designed to resemble a virtual small town with a simple street layout consisting of four vertical and four horizontal roads, as shown in Figure 1. We designed a two-step navigation task to evaluate each locomotion method within a task-oriented context. We first instructed participants to locate a bank in the virtual town. Each participant started at the same position, marked by the blue “X” in Figure 1, and completed the task by finding the target locations at identical coordinates. Since the main goal was to compare navigation methods rather than participants’ ability to find targets, we made the task intuitive by placing a prominent pink hologram in front of the bank, making it easily identifiable. Upon reaching the bank, the first hologram disappeared, and a second one appeared at the town’s exit as a second target. Participants were then required to find and reach this exit point to complete the navigation task. It is important to note that in the LLM-based mode, participants could only travel short visible distances of 50 meters. Therefore, they could not directly say ‘go to the bank.’
@@ 2.2.4 Procedure
We first welcomed the participants and asked them to provide written consent to participate in the experiment. Next, we explained the (randomly) assigned locomotion condition to them in detail. Participants then wore the HMD, and the session began with a demo scene designed to help them become familiar with the locomotion technique and VR. The demo scene used the same virtual environment, starting from the same initial position, and included a single target located at the first corner, which was easily visible and close to the starting point. Eye-tracking calibration was also performed during this phase. Once participants felt comfortable about the VR and controls, following the demo, the main experiment commenced. As instructed, participants were supposed to navigate the virtual environment and reach two holograms placed at specific locations. After completing the task, participants completed questionnaires, including the NASA-TLX, SUS, IPQ, CSQ-VR, and a demographic survey. Later, they received €7.5 for the 30-minute experimental session.
@@ 2.2.5 Measurements
To evaluate the locomotion techniques, we conducted an eye-tracking analysis, as eye tracking has been widely used to assess user engagement, attention, behavior, and cognitive processes during interaction [38, 80]. We classified samples as either fixations or saccades from the raw eye-tracking data. Fixations represent moments when the eyes remain steadily focused on a specific point, while saccades are rapid eye movements between fixation points. We analyzed several features to understand users’ visual attention and cognitive states. The fixation rate, which reflects the frequency of attentional shifts, is commonly used to assess how users allocate their attention during a task [28]. Mean fixation duration, which represents the length of time users dwell on specific points, has been linked to deeper cognitive processing and enhanced information encoding [43]. The mean saccade duration captures the speed of gaze transitions, indicating the efficiency of visual search behavior. Pupil diameter provides insights into the mental effort and arousal, correlating with the cognitive workload, fatigue, and engagement [6, 40]. Additionally, we identified blinks and extracted both blink frequency and duration, as both are associated with cognitive load and visual fatigue [7].
@@ 2.2.7 Analysis
We conducted a series of statistical analyses on both questionnaire responses (e.g., SUS, NASA-TLX) and eye-tracking measures to examine differences across experimental conditions. For each dependent variable, we performed a one-way Analysis of Variance (ANOVA) to evaluate the effect of the locomotion technique (teleportation, LLM-driven locomotion, and voice-based steering). When the ANOVA indicated a statistically significant effect, we used Tukey’s Honest Significant Difference (HSD) post-hoc test to determine which pairs of conditions differed significantly, thereby controlling the family-wise error rate across all comparisons.
Before conducting ANOVA, we checked the assumptions of normality with the Shapiro–Wilk test and homogeneity of variance with Levene’s test. In cases where these assumptions were violated, we conducted the non-parametric Kruskal-Wallis H test as an alternative to ANOVA. We conducted all statistical tests using . Where relevant, we also report the effect sizes (e.g., ) to support the interpretation of findings.
@@ 2.2.8 Model Building
We conducted detailed analyses of eye-tracking data to investigate how different locomotion techniques influence user behavior. These analyses aim to reveal potential cognitive and attentional differences among participants using teleportation, voice-based steering locomotion, and LLM-driven locomotion techniques. Additionally, we trained classification models to predict the locomotion condition based on eye-tracking features. This exploratory analysis helps identify which features are most affected by the locomotion technique and which features are the most predictive or distinctive in differentiating between the locomotion techniques. We provide the full list of features used in Table 2.
@@
To this end, we evaluated five classifiers: Random Forest, LightGBM, k-Nearest Neighbor (k-NN), Support Vector Machine (SVM), and Gradient Boosting. The classification was performed on features from 20-second windows, by splitting 80% of the data for training and 20% for testing. We applied hyperparameter optimization using a stratified 5-fold cross-validation to ensure robust model performance. To understand the models’ decision-making processes, we applied SHAP (SHapley Additive exPlanations) [50] to interpret feature importance in a transparent and model-agnostic fashion. SHAP values provide a detailed attribution of feature contributions toward classification outcomes, helping us identify which aspects of gaze behavior, such as fixation duration, saccade amplitude, or pupil diameter, are most influenced by the underlying locomotion technique, which can also help personalize VR.
@@ 3.2 Eye-Tracking Analysis
To assess cognitive engagement under different locomotion techniques, we analyzed the fixation rate, defined as the number of fixations per second, as shown in Figure 6. ANOVA revealed a statistically significant effect of locomotion type on fixation rate, , , with a large effect size (), indicating that the type of locomotion significantly influenced visual attention patterns. Additionally, the statistics indicated that the voice-based steering condition resulted in the highest fixation rate (, ), followed by teleportation (, ), and voice LLM (, ). Post-hoc comparisons using Tukey’s Honest Significant Difference (HSD) test showed that both teleportation and voice LLM conditions had significantly lower fixation rates compared to the voice-based steering condition ( for both comparisons). However, no statistically significant difference was found between the teleportation and voice LLM conditions (), indicating comparable visual engagement.
@@
Mean saccade durations for each condition are presented in Figure 6, and all values are reported in milliseconds. ANOVA revealed a statistically significant effect of locomotion type on mean saccade duration, , , with a large effect size (). Descriptive statistics showed the highest mean saccade duration in voice-based steering (, ms), followed by teleportation (, ms), and the lowest in the LLM-driven condition (, ms). Post-hoc tests showed mean saccade duration was significantly lower in the LLM-driven condition than in teleportation () and voice-based steering (), with no difference between the latter two ().